OpenStudy (anonymous):

What can I compare the integral of 1/x+ srqtx from 1 to infinity to in order to find whether it converges?

OpenStudy (anonymous):

you mean what infinite sum?

OpenStudy (anonymous):

i guess it would be \[\sum_{k=1}^{\infty}\frac{1}{n+\sqrt{n}}\]